Which of these best denotes total investment in GDP?

C+Y+G

S-C-G

C+G+NX

Y-C-G

The government raised $12m in taxes and spent $10m on every budget item, consumers spent roughly $30m on goods and services, and total GDP was $60m. What would public saving be? (Please show work)

-$20m

-$22m

-$18m

- $2m
Answer #1

Q1

Answer

the equilibrium equation is

Y=C+I+G

I=Y-C-G

Option 4

----------

Q2

ANswer

Option 4

Public saving =T-G=taxes - government spending

=12-10

=$2 million

the public saving is $2 million